OpenStudy (anonymous):

Describe the difference between the graph of: y - 3 = -tan( -2x + 180°) and y = tan(x)

OpenStudy (anonymous):

The general form of the tangent function could be written as \[y=a\tan \big(b(x-c)\big)+d\] where \(a\) is a scaling factor, \(b\) is related to the period of the function (with \(p=\dfrac{\pi}{b}\)), \(c\) is a horizontal shift of \(c\) units left (if positive) or right (if negative), and \(d\) is a vertical shift of \(d\) units up (if positive) or down (if negative). Try writing \(y-3=-\tan\left(-2x+180^\circ \right)\) in this general form.

OpenStudy (anonymous):

ok so would say the differences then?

OpenStudy (anonymous):

how would **

OpenStudy (anonymous):

The general form of your given function is \[y=-\tan\big(-2(x-90^\circ)\big)+3\] So you have \(a=-1\), \(b=-2\), \(c=90^\circ\), and \(d=3\). \(a\) is the scaling factor. In this case, it's negative, so the change made to the graph of \(\tan x\) is a reflection across the x-axis.